# GTAvatar: Bridging Gaussian Splatting and Texture Mapping for Relightable and Editable Gaussian Avatars

### v1

@@ 3.4 UV-Mapping for Textured Gaussians
Our method extends Gaussian Splatting by replacing the color of Gaussians with texture patches. Contrary to existing works that enhance the representation power of Gaussians with small per-primitive textures [gstex, BBSplat, texturedgaussiansenhanced3d], we seek to embed Gaussians into a single texture space with a semantic mapping predefined on the template mesh. To achieve this, we map the tangent plane of each splat to a continuous patch in the mesh’s UV space. We modify the Gaussian Splatting rasterizer to retrieve colors by sampling a texture at the mapped UV coordinates efficiently with bilinear filtering. This texture is learned alongside the Gaussian parameters.
@@
where $J_{\text{v}}^{\dagger}$ is the pseudo-inverse of $J_{\text{v}}$ and $\text{uv}_{0}$ the UV coordinates at the center of the splat, computed using the Gaussian’s barycentric coordinates directly. We extend this to Gaussians that are not aligned with their triangle by noting that $J_{\text{v}}J_{\text{v}}^{\dagger}$ is the orthogonal projector onto the column space of $J_{\text{v}}$ (the triangle plane) and $x\mapsto J_{\text{v}}^{\dagger}(x-\text{V}_{A})$ yields the barycentric coordinates of the projection. Thus, $\mathcal{M}$ maps $(s,t)$ to the UV coordinates of this projection (the $J_{\text{v}}^{\dagger}\text{V}_{A}$ term is accounted for in $\text{uv}_{0}$).
This lightweight mapping only requires a single matrix multiplication at each ray-splat intersection during rendering. The largest computational costs are the pseudo-inversion of $J_{\text{v}}^{\dagger}$ for each triangle and the bilinear texture filtering; the latter is a well-understood cost of graphics pipelines that can be mitigated with hardware acceleration. Notice that $J_{\text{v}}$ – and its pseudo-inverse $J_{\text{v}}^{\dagger}$ – are expression-dependent but do not need to be updated for viewpoint changes. See Section 5.5 for more details on rendering speed and a comparison with a naive projection baseline.
@@ 3.5.1 Real-time Physically-Based Rendering
We adopt the real-time shading model of [karis2013real], which we briefly summarize in this section. The rendering equation expresses the outgoing light $L_{0}$ leaving surface point $x$ in the camera direction $\mathbf{\omega_{0}}$ by integrating the light reaching it from all directions:
@@
where $f_{r}$ is the bidirectional reflectance distribution function of the material (BRDF). We decompose the BRDF as a diffuse Lambertian term independent of view direction:
@@
and a view-dependent specular term, calculated using the split-sum approximation [karis2013real]:
@@
Similarly to previous relightable head avatars [nvdiffrec, flare, HRAvatar], we adopt the Cook-Torrance microfacet reflectance model [cook1982reflectance], which parameterizes the specular BDRF with surface roughness $r$ and specular reflectance at normal incidence $f_{0}$. In practice, the second integral incorporates parts of $f_{r,\text{spec}}$ that are not dependent on $\mathbf{\omega_{0}}$, and is stored as a pre-filtered environment map $\mathcal{L}(\mathbf{\omega_{r}},r)$ with mip levels sampled by spatially-varying roughness $r$, where $\mathbf{\omega_{r}}$ is the reflection of $\mathbf{\omega_{0}}$ onto the normal. We learn this function as a RGB cubemap that represents the lighting of the scene. The first integral is a function of material properties and viewing angle, independent of lighting. It is pre-computed in a look-up table; for conciseness, we refer the reader to [karis2013real] for details and denote this reflectance term as $R(f_{0},r,\mathbf{\omega_{0}}\cdot\mathbf{n})$. Note that due to properties of the BRDF omitted here, the integral in Eq. 3 can be computed as $\mathcal{L}(\mathbf{n},0)$. Hence the final color is:
@@ 3.5.2 Normal Mapping
[zielonka2022insta] INSTA $\times$ $\times$ 27.85 0.9110 0.1047 25.03 0.8475 0.1614
[pointavatar] Point-avatar $\times$ $\times$ 26.84 0.8970 0.0926 25.14 0.8385 0.1278
[splattingavatar] Splatting-avatar $\times$ $\times$ 28.71 0.9271 0.0862 26.66 0.8611 0.1351
[flashavatar] Flash-avatar $\times$ $\times$ 29.13 0.9255 0.0719 27.58 0.8664 0.1095
[gaussianblendshapes] GBS $\times$ $\times$ 29.64 0.9394 0.0823 27.81 0.8915 0.1297
[flare] FLARE $\times$ 26.80 0.9063 0.0816 25.55 0.8479 0.1183
@@
Figure 6: Reconstruction quality on unseen frames averaged over 10 videos of the INSTA dataset, compared with HRAvatar [HRAvatar] (LPIPS, lower is better). Densification and pruning of Gaussians are disabled in these experiments for manual control of splat count.
@@ 4 Training
Our model is trained end-to-end from a monocular RGB video. Following previous work [pointavatar, flare, HRAvatar], the input video is preprocessed prior to training to extract alpha masks using a matting method [robustvideomatting] and initial FLAME parameters using an off-the-shelf facial tracker [SMIRK]. During training, we simultaneously optimize the parameters of our 2D Gaussians (barycentric coordinates, relative rotation, scale, displacement and opacity), the textures (albedo, roughness, specular reflectance and normals) and the scene’s lighting environment map (a $6\times 32^{2}\times 3$ RGB cubemap). Gaussians undergo densification and pruning as in other Gaussian-based methods [3DGS, 2DGS, gaussianavatars, HRAvatar]. We also optimize the FLAME model for higher fidelity and fine-tune the expression encoder from the FLAME tracker [SMIRK] while continuously predicting updated expression parameters during training, following recent work [HRAvatar]. We train for 15 epochs using Adam optimizer [adam], requiring one to two hours on an NVIDIA RTX A5000 GPU. Learning rates are given in Table 3(a).
@@
Values used for the weights referenced in this section are provided in Table 3(b).
@@ 4.2 Priors for Physical Disentanglement
First, we supervise the rendered albedo $\hat{I}_{\rho}$ with a diffusion-based prior $I_{\rho}$ [chen2024intrinsicanything] precomputed for every third frame of the video, as recent work has proven this effective for disentangling albedo and lighting [HRAvatar].
@@ 4.3 UV-Mapping Regularization
UV Distortion Loss. Inspired by the depth distortion loss in 2DGS [2DGS], we propose a UV distortion loss to concentrate the texture coordinates of all ray-splat intersections contributing to a single ray. For a given pixel ray, we minimize the weighted pairwise distance between the UV coordinates of all ray-splat intersections. Let $uv_{i}$ be the $(u,v)$ coordinate derived from the $i$-th splat intersection along a ray. The loss then writes:
@@
where $\mathcal{T}_{\text{init}}$ is the initial value for that texture – $[0,0,0]$ for albedo, $0.5$ for roughness, $0.05$ for specular reflectance and $[=0,0,1]$ for normals.
@@
Together, $\mathcal{L}_{\text{uv}}=\lambda_{\text{uv\_dist}}\mathcal{L}_{\text{uv\_dist}}+\lambda_{\text{boundary}}\mathcal{L}_{\text{uv\_boundary}}+\lambda_{\text{stat\_albedo}}\mathcal{L}_{\text{stat\_albedo}}$. These terms are critical for producing high-quality, editable textures as shown in Figure 9.
@@ 4.4 Geometric and Gaussian Regularization
Surface Normal Consistency. To ensure that our 2D splats align locally with the overall reconstructed surface, we adopt the normal consistency loss from 2DGS [2DGS]. This loss aligns the normal of each individual splat $\mathbf{n}_{i}$ with the macro-scale surface normal $\mathbf{N}$, derived from the screen-space gradients of the rendered depth map. The loss for each ray is expressed as:
@@
Reconstruction Textured edit Texture
@@ 5 Results
Dataset. We evaluate various aspects of our method on two common datasets for monocular avatar reconstruction: INSTA [zielonka2022insta] and HDTF [zhang2021flow]. Both provide 2-3 minute $512\times 512$ resolution talking head videos. Our experimental setting aligns with that of our main baseline, HRAvatar [HRAvatar]: we use 10 subjects from the INSTA dataset, and 8 from HDTF. For self-reenactment evaluation, the last 350 frames are left out of training for the former, and the last 500 frames for the latter.
@@ 5.1 Self-reenactment
Figure 10: Examples of reconstructed images and novel viewpoints.
@@
Figure 11: Visual comparison of our results with FATE [fate]. While both approaches provide texture editing features, our UV mapping technique yields significantly sharper renders with high-frequency details. Numbers indicate the respective number of splats.
@@
Figure 12: Visual comparison of our method with HRAvatar [HRAvatar] and FLARE [flare] for reconstruction and environment map relighting.
In this section, we validate the quality of our reconstructions in the self-reenactment setting. Unseen frames from the video are tracked to recover FLAME parameters using the tuned encoder, then reconstructed. In Table 1 we report average PSNR, SSIM and VGG-based LPIPS metrics for the INSTA and HDTF datasets, compared to state-of-the-art monocular avatar reconstruction methods as evaluated by the authors of HRAvatar [HRAvatar]. Our method displays state of the art performance, comparably with HRAvatar, outperforming all baselines. Qualitative reconstruction examples are shown in Figures 4 and 12. Figure 10 showcases our method’s ability to render realistic avatars from novel viewpoints.
Moreover, we assess the effect of primitive count and texture resolution in Figure 6. Average LPIPS on the INSTA dataset is reported for various primitive counts and texture resolutions. For fairness to HRAvatar which does not have textures, we also compare with equal parameter counts. These results underline the efficiency of our method, enabling on-par reconstruction quality with a significantly smaller model size. The added representation power of textures reduces the number of primitives required through a compact representation of normal and appearance attributes enabled by the UV-space locality of adjacent splats.
@@ 5.2 Texture Editing
In Figure 9, we perform an ablation study of several aspects of our method aimed at improving the texture mapping and the quality of reconstructed textures. First, we validate that the UV distortion loss enhances the sharpness of our render. While the reconstruction converges well without the regularization, mapping a texture with high-frequency details reveals some blurriness. Next, we show that the statistical albedo regularization is able to fill-in the holes in the textures, remove distracting artifacts and better preserving the structure of the texture space. Finally, we set our Jacobians (see Section 3.4) to zero, which equates to using a single UV coordinate per splat. This per-primitive mapping yields a discontinuous texture with only a sparse set of texels being used (note that we disable the statistical albedo regularization to better visualize those texels), and cannot be edited at high resolution unless the number of primitives is impractically high.
FATE [fate] allows texture-based editing of Gaussian Avatars. However, their method relies on per-primitive color, which severely limits the quality of rendering for high-frequency textures. Figure 11 compares renders with the same texture for FATE and GTAvatar. Our method outperforms FATE even while increasing the number of primitives FATE uses well beyond its default.
@@ 5.3 Normal mapping
In Figure 8, we demonstrate the impact of our normal mapping on the sharpness of rasterized normals and the quality of relit rendering. Furthermore, Figure 5 shows how handcrafted normal maps substantially enhance the realism for material editing under dynamic lighting.
@@ 5.5 Rendering time
Sampling a texture at every ray-splat intersection comes with a computational cost. To minimize this cost, our implementation uses hardware acceleration at test time for more efficient access via CUDA texture objects. In Table 2, we compare rendering speed with other methods, and with variants of our method without hardware acceleration, with more Gaussians and with naive projection instead of the fast UV mapping described in Section 3.4. In the latter, we explicitly perform orthogonal projection to the triangle plane and calculate barycentric coordinates at ray-splat intersection. We report FPS for both static and dynamic geometry, as we found FLAME deformations and, in our case, updating the Jacobians, to be a significant contributor to rendering time. Note that the static geometry case still enables camera movement and relighting. While our method performs slower than the HRAvatar baseline for a given number of primitives, ours requires less (as shown in Figure 6(a)), yielding on par rendering time for comparable or better reconstruction quality. The final trained avatar can be rendered at more than 170 FPS with static geometry, or 80 FPS with dynamic geometry on a RTX A5000.
@@
Table 2: Comparison of rendering speed at inference for static and dynamic geometry on a NVIDIA RTX A5000 GPU, and average LPIPS unseen frames of the INSTA dataset. Numbers next to method names indicate the number of Gaussians used, when applicable (densification and pruning are disabled for manual control). Despite the additional overhead of texture sampling, our method achieves competitive speed by reducing the number of Gaussians required to achieve a photorealistic reconstruction.
@@ 7 Conclusion
As part of future work, several avenues remain to further improve our representation. Reducing aliasing artifacts is a key challenge: incorporating anti-aliasing strategies, either through trilinear texture filtering in the UV domain or directly within the 2DGS splatting stage [AA2DGS] could enhance visual smoothness and stability of avatars and edits. In addition, extending this representation to the better-constrained multi-view setting could enable further uses. Such extensions would make our method even more suitable for production-grade applications.

### v2

@@ 3.4 UV-Mapping for Textured Gaussians
Our method extends Gaussian Splatting by replacing the color of Gaussians with texture patches. Contrary to existing works that enhance the representation power of Gaussians with small per-primitive textures [64, 76, 9], we seek to embed Gaussians into a single texture space with a semantic mapping predefined on the template mesh. To achieve this, we map the tangent plane of each splat to a continuous patch in the mesh’s UV space. We modify the Gaussian Splatting rasterizer to retrieve colors by sampling a texture at the mapped UV coordinates efficiently with bilinear filtering. This texture is learned alongside the Gaussian parameters.
@@
where $J_{\text{v}}^{\dagger}$ is the pseudo-inverse of $J_{\text{v}}$ and $\text{uv}_{0}$ the UV coordinates at the center of the splat, computed using the Gaussian’s barycentric coordinates directly. We extend this to Gaussians that are not aligned with their triangle by noting that $J_{\text{v}}J_{\text{v}}^{\dagger}$ is the orthogonal projector onto the column space of $J_{\text{v}}$ (the triangle plane) and $x\mapsto J_{\text{v}}^{\dagger}(x-\text{V}_{A})$ yields the barycentric coordinates of the projection. Thus, $\mathcal{M}$ maps $(s,t)$ to the UV coordinates of this projection (the $J_{\text{v}}^{\dagger}\text{V}_{A}$ term is accounted for in $\text{uv}_{0}$). The pseudo-inversion is implemented by solving a least-squares problem, which does not back-propagate gradients to the template vertices.
This lightweight mapping only requires a single matrix multiplication at each ray-splat intersection during rendering. The largest computational costs are the pseudo-inversion of $J_{\text{v}}^{\dagger}$ for each triangle and the bilinear texture filtering; the latter is a well-understood cost of graphics pipelines that can be mitigated with hardware acceleration. Notice that $J_{\text{v}}$ – and its pseudo-inverse $J_{\text{v}}^{\dagger}$ – are mesh and expression-dependent but do not need to be updated for viewpoint changes. See Section 5.5 for more details on rendering speed and a comparison with a naive projection baseline.
@@ 3.5.1 Real-time Physically-Based Rendering
We adopt the real-time shading model of [34], which we briefly summarize in this section. The rendering equation expresses the outgoing light $L_{0}$ leaving surface point $x$ in the camera direction $\mathbf{\omega_{0}}$ by integrating the light reaching it from all directions:
@@
where $f_{r}$ is the bidirectional reflectance distribution function of the material (BRDF). We decompose the BRDF as a diffuse Lambertian term independent of incoming light direction:
@@
and a view-dependent specular term, calculated using the split-sum approximation [34]:
@@
Similarly to previous relightable head avatars [52, 4, 101], we adopt the Cook-Torrance microfacet reflectance model [15], which parameterizes the specular BDRF with surface roughness $r$ and specular reflectance at normal incidence $f_{0}$. In practice, the second integral incorporates parts of $f_{r,\text{spec}}$ that are not dependent on $\mathbf{\omega_{0}}$, and is stored as a pre-filtered environment map $\mathcal{L}(\mathbf{\omega_{r}},r)$ with mip levels sampled by spatially-varying roughness $r$, where $\mathbf{\omega_{r}}$ is the reflection of $\mathbf{\omega_{0}}$ onto the normal. We learn this function as a RGB cubemap that represents the lighting of the scene. The first integral is only dependent on material properties and viewing angle and independent of lighting. It is pre-computed in a look-up table; for conciseness, we refer the reader to [34] for details and denote this reflectance term as $R(f_{0},r,\mathbf{\omega_{0}}\cdot\mathbf{n})$. Note that due to properties of the BRDF omitted here, the integral in Eq. 3 can be computed as $\mathcal{L}(\mathbf{n},0)$. Hence the final color is:
@@ 3.5.2 Normal Mapping
[109] INSTA $\times$ $\times$ 27.85 0.9110 0.1047 25.03 0.8475 0.1614
[106] Point-avatar $\times$ $\times$ 26.84 0.8970 0.0926 25.14 0.8385 0.1278
[69] Splatting-avatar $\times$ $\times$ 28.71 0.9271 0.0862 26.66 0.8611 0.1351
[86] Flash-avatar $\times$ $\times$ 29.13 0.9255 0.0719 27.58 0.8664 0.1095
[46] GBS $\times$ $\times$ 29.64 0.9394 0.0823 27.81 0.8915 0.1297
[4] FLARE $\times$ 26.80 0.9063 0.0816 25.55 0.8479 0.1183
@@
Figure 6: Reconstruction quality on unseen frames averaged over 10 videos of the INSTA dataset, compared with HRAvatar [101] (LPIPS, lower is better). Densification and pruning of Gaussians are disabled in these experiments for manual control of splat count.
@@ 4 Training
Our model is trained end-to-end from a monocular RGB video. Following previous work [106, 4, 101], the input video is preprocessed prior to training to extract alpha masks using a matting method [45] and initial FLAME parameters using an off-the-shelf facial tracker [63]. During training, we simultaneously optimize the parameters of our 2D Gaussians (barycentric coordinates, relative rotation, scale, displacement and opacity), the textures (albedo, roughness, specular reflectance and normals) and the scene’s lighting environment map (a $6\times 32^{2}\times 3$ RGB cubemap). Gaussians undergo densification and pruning as in other Gaussian-based methods [35, 31, 60, 101]. We also optimize the FLAME model (template mesh and deformation basis) for higher fidelity and fine-tune the expression encoder from the FLAME tracker [63] while continuously predicting updated expression parameters during training, following recent work [101]. We train for 15 epochs using Adam optimizer [38], requiring one to two hours on an NVIDIA RTX A5000 GPU. Learning rates are given in Table 3a.
@@
Values used for the weights referenced in this section are provided in Table 3b.
@@ 4.2 Priors for Physical Disentanglement
First, we supervise the rendered albedo $\hat{I}_{\rho}$ with a diffusion-based prior $I_{\rho}$ [84] precomputed for every third frame of the video, as recent work has proven this effective for disentangling albedo and lighting [101].
@@ 4.3 UV-Mapping Regularization
UV Distortion Loss. Inspired by the depth distortion loss in 2DGS [31], we propose a UV distortion loss to concentrate the texture coordinates of all ray-splat intersections contributing to a single ray. For a given pixel ray, we minimize the weighted pairwise distance between the UV coordinates of all ray-splat intersections. Let $uv_{i}$ be the $(u,v)$ coordinate derived from the $i$-th splat intersection along a ray. The loss then writes:
@@
where $\mathcal{T}_{\text{init}}$ is the initial value for that texture – $[0,0,0]$ for albedo, $0.5$ for roughness, $0.05$ for specular reflectance and $[0,0,1]$ for normals.
@@
Together, $\mathcal{L}_{\text{uv}}=\lambda_{\text{uv\_dist}}\mathcal{L}_{\text{uv\_dist}}+\lambda_{\text{boundary}}\mathcal{L}_{\text{uv\_boundary}}+\lambda_{\text{stat\_albedo}}\mathcal{L}_{\text{stat\_albedo}}$. These terms are critical for producing high-quality editable textures as shown in Figure 9.
@@ 4.4 Geometric and Gaussian Regularization
Surface Normal Consistency. To ensure that our 2D splats align locally with the overall reconstructed surface, we adopt the normal consistency loss from 2DGS [31]. This loss aligns the normal of each individual splat $\mathbf{n}_{i}$ with the macro-scale surface normal $\mathbf{N}$, derived from the screen-space gradients of the rendered depth map. The loss for each ray is expressed as:
@@
Reconstruction Texture edit Rec. texture
@@ 5 Results
Dataset. We evaluate various aspects of our method on two common datasets for monocular avatar reconstruction: INSTA [109] and HDTF [105]. Both provide 2-3 minute $512\times 512$ resolution talking head videos. Our experimental setting aligns with that of our main baseline, HRAvatar [101]: we use 10 subjects from the INSTA dataset, and 8 from HDTF. For self-reenactment evaluation, the last 350 frames are left out of training for the former, and the last 500 frames for the latter.
@@ 5.1 Self-reenactment
Figure 10: Examples of reconstructions and novel viewpoints.
@@
Figure 11: Visual comparison of our results with FATE [102]. While both approaches provide texture editing features, our UV mapping technique yields sharper renders with high-frequency details. Numbers indicate the respective number of splats.
@@
Figure 12: Visual comparison of our method with HRAvatar [101] and FLARE [4] for reconstruction and environment map relighting.
In this section, we validate the quality of our reconstructions in the self-reenactment setting. Unseen frames from the video are tracked to recover FLAME parameters using the tuned encoder, then reconstructed. In Table 1 we report average PSNR, SSIM and VGG-based LPIPS metrics for the INSTA and HDTF datasets, compared to state-of-the-art monocular avatar reconstruction methods as reported by the authors of HRAvatar [101]. Our method displays state of the art performance, comparably with HRAvatar, outperforming all baselines. Qualitative reconstruction examples are shown in Figures 4 and 12. Figure 10 showcases our method’s ability to render realistic avatars from novel viewpoints.
Moreover, we assess the effect of primitive count and texture resolution in Figure 6. Average LPIPS on the INSTA dataset is reported for various primitive counts and texture resolutions. For fair comparison with HRAvatar which does not have textures, we also compare with equal parameter counts. These results underline the efficiency of our method, enabling on-par reconstruction quality with a significantly smaller model size. The added representation power of textures reduces the number of primitives required through a compact representation of normal and appearance attributes enabled by the UV-space locality of adjacent splats.
@@ 5.2 Texture Editing
In Figure 9, we perform an ablation study of several aspects of our method aimed at improving the texture mapping and the quality of reconstructed textures. First, we validate that the UV distortion loss enhances the sharpness of our render. While the reconstruction converges well without the regularization, mapping a texture with high-frequency details reveals some blurriness. Next, we show that the statistical albedo regularization is able to fill-in the holes in the textures, remove distracting artifacts and better preserve the structure of the texture space. Finally, we set our Jacobians (see Section 3.4) to zero, which equates to using a single UV coordinate per splat. This per-primitive mapping yields a discontinuous texture with only a sparse set of texels being used (note that we disable the statistical albedo regularization to better visualize those texels), and cannot be edited at high resolution unless the number of primitives is impractically high.
FATE [102] allows texture-based editing of Gaussian Avatars. However, their method relies on per-primitive color, which severely limits the quality of rendering for high-frequency textures. Figure 11 compares renders with the same texture for FATE and GTAvatar. Our method outperforms FATE even while increasing the number of primitives FATE uses well beyond its default.
@@ 5.3 Normal mapping
In Figure 8, we demonstrate the impact of our normal mapping on the sharpness of rasterized normals and the quality of relit renders. Furthermore, Figure 5 shows how handcrafted normal maps substantially improve the realism for material editing under dynamic lighting.
@@ 5.5 Rendering time
Sampling a texture at every ray-splat intersection comes with a computational cost. To minimize this cost, our implementation uses hardware acceleration at test time for more efficient access via CUDA texture objects. In Table 2, we compare rendering speed with other methods, and with variants of our method without hardware acceleration, with more Gaussians and with naive projection instead of the fast UV mapping described in Section 3.4. In the latter, we explicitly perform orthogonal projection to the triangle plane and calculate barycentric coordinates at ray-splat intersection. We report FPS for both static and dynamic geometry, as we found FLAME deformations and, in our case, updating the Jacobians, to be a significant contributor to rendering time. Note that the static geometry case still enables camera movement and relighting. While our method performs slower than the HRAvatar baseline for a given number of primitives, ours requires less (as shown in Figure 6a), yielding on par rendering time for comparable or better reconstruction quality. The final trained avatar can be rendered at more than 170 FPS with static geometry, or 80 FPS with dynamic geometry on a RTX A5000.
@@
[…]
Table 2: Comparison of rendering speed at inference for static and dynamic geometry on a NVIDIA RTX A5000 GPU, and average LPIPS for unseen frames of the INSTA dataset. Numbers next to method names indicate the number of Gaussians used, when applicable (densification and pruning are disabled for manual control). Despite the additional overhead of texture sampling, our method achieves competitive speed by reducing the number of Gaussians required to achieve a photorealistic reconstruction.
@@ 7 Conclusion
Several limitations remain. The use of image-based lighting to represent scene illumination enables intuitive manipulation of the environment, but comes with shortcomings. Most importantly, this approach can only model light at infinity and is unable to represent point lights. Furthermore, effects such as indirect illumination and shadows are not represented, and will thus be baked into other properties, yielding inaccurate physical decomposition.
As part of future work, several avenues remain to further improve our representation. The texture-mapping contribution could be integrated in a Gaussian rasterizer that enables more advanced lighting representations, e.g. through ray-tracing [50, 87, 27, 74]. Reducing aliasing artifacts is also a key challenge: incorporating anti-aliasing strategies, either through trilinear texture filtering in the UV domain or directly within the 2DGS splatting stage [95] could enhance visual smoothness and stability of avatars and edits. In addition, extending this representation to the better-constrained multi-view setting could enable further uses. Such extensions would make our method even more suitable for production-grade applications.